Georgian Mathematical Journal Volume 9 (2002), Number 1, 197–

A SEMIMARTINGALE BELLMAN EQUATION AND THE

VARIANCE-OPTIMAL MARTINGALE MEASURE. CORRECTION

M. MANIA AND R. TEVZADZE

Abstract. Corrections are given for [1], most of which consist in replacing “ess sup” by “ess inf ” in all the exposed formulas.

2000 Mathematics Subject Classification: 91B28, 60H30, 90C39. Key words and phrases: Backward semimartingale equation, incomplete markets, contingent claim pricing, variance-optimal martingale measure, the Bellman equation.

Unfortunately, [1] contains some distortions of the meaning most of which can be corrected as follows: replace throughout the paper “ess sup” by “ess inf” in all the exposed formulas; delete the second summand of μ(t, x) in the third line of formula (4.17).

References

1. M. Mania and R. Tevzadze, A semimartingale Bellman equation and the varianceoptimal martingale measure. Georgian Math. J. 7(2000), No. 4. 765–792.

ISSN 1072-947X / $8.00 / \begin{matrix}c \\ ◯ \end{matrix} Heldermann Verlag www.heldermann.de